OpenStudy (anonymous):

How do I solve 2x^4-32=0

OpenStudy (paxpolaris):

\[2x^4-32=0 \]\[\implies2x^4=32\]

OpenStudy (paxpolaris):

\[\implies x^4=16\]

OpenStudy (paxpolaris):

real solutions: 2, -2

OpenStudy (anonymous):

Are there any complex solutions? My book is giving me \[\pm2\] and \[\pm2i\]

OpenStudy (paxpolaris):

that is correct: \[x^4=16\] \(\implies x^2=4\), or \(x^2=-4\)

OpenStudy (anonymous):

Alright, got it . Thanks